# Stability Regularized Cross-Validation

### v1

@@ Abstract
We revisit the problem of ensuring strong test-set performance via cross-validation. Motivated by the generalization theory literature, we propose a nested k-fold cross-validation scheme that selects hyperparameters by minimizing a weighted sum of the usual cross-validation metric and an empirical model-stability measure. The weight on the stability term is itself chosen via a nested cross-validation procedure. This reduces the risk of strong validation set performance and poor test set performance due to instability. We benchmark our procedure on a suite of $13$ real-world UCI datasets, and find that, compared to $k$-fold cross-validation over the same hyperparameters, it improves the out-of-sample MSE for sparse ridge regression and CART by $4\%$ on average, but has no impact on XGBoost. This suggests that for interpretable and unstable models, such as sparse regression and CART, our approach is a viable and computationally affordable method for improving test-set performance.
@@ 1 Introduction
A central problem in machine learning and data-driven optimization involves constructing models that reliably generalize well to unseen data. One of the most popular approaches is cross-validation as introduced by [53, 18], which selects hyperparameters that perform well on a cross-validation set as a proxy for strong test-set performance. Indeed, this model selection pipeline is advocated by most machine learning textbooks [e.g. 28, 25]. Moreover, in the statistical learning literature, there is a broad set of conditions under which the cross-validation loss (possibly with mild corrections) is a good estimator of the test-set error [41, 32, 29, 40], especially when the sample size is large [34, 11].
@@
At the same time, both theory [e.g. 48, 22] and experiments [e.g. 45, 44, 42, 52, 3] have documented an adaptivity gap between validation and test-set performance (although not always observed [46, 3]), especially in settings with limited data. Concretely, a positive adaptivity gap arises when the validation error is systematically lower than the test set error of a machine learning model.
@@
One explanation for adaptivity gaps is as follows: validation sets give approximately unbiased estimators of test set performance for a fixed combination of hyperparameters. However, validation scores are random variables and subject to some variance. Therefore, the act of optimizing the validation set error risks selecting hyperparameter combinations that disappoint out of sample. In the extreme case with many hyperparameters relative to the number of samples, the act of optimizing the (cross) validation error can be viewed as training on the (cross) validation set. This phenomenon is well documented in different parts of the statistics and optimization literature, where it is variously called “post-decision surprise”, “out-of-sample disappointment”, “researcher degree of freedom” or “the optimizer’s curse” [27, 38, 39, 51, 50, 56]. This raises the question: is it possible to improve the performance of (cross) validation by combating the adaptivity gap, possibly by selecting models with a higher validation error and a lower score according to a second metric.
@@
In this work, we propose a strategy for mitigating out-of-sample disappointment and show that it sometimes improves cross-validation’s performance. Specifically, we propose selecting models according to a weighted sum of their cross-validation error and their empirical hypothesis stability (see Section 2.2). This is motivated by the observation that both the cross-validation error and the hypothesis stability appear in generalization bounds on the test-set error; thus, minimizing the cross-validation error alone may be vulnerable to selecting high-variance models that perform poorly out-of-sample.
@@
Our main contributions are threefold. First, we extend a generalization bound on the test set error due to [11] from leave-one-out to $k$-fold cross-validation. This generalization bound takes the form of the cross-validation error plus a term related to a model’s hypothesis stability. Second, motivated by this (often conservative) bound, we propose regularizing cross-validation by selecting models that minimize a weighted sum of a validation metric and the hypothesis stability, rather than the validation score alone, to mitigate out-of-sample disappointment without being overly conservative. Indeed, models with a low cross-validation error that are stable generalize better than models with a low cross-validation error that are unstable. Moreover, to select the weight in this scheme, we embed the entire scheme within a nested cross-validation procedure. Finally, we empirically investigate our proposal using sparse ridge regression, CART, and XGBoost models, and find that it improves the out-of-sample performance of sparse ridge regression and CART by $4\%$ on average but has no impact on XGBoost, likely because XGBoost is a more stable learner.
@@ 1.1 Motivating Example: Poor Performance of Cross-Validation for Sparse Linear Regression
We illustrate the pitfalls of cross-validation in a sparse ridge regression setting, as studied by [7, 31, 35]. Suppose we wish to recover a $\tau_{\text{true}}=5$ sparse regressor which is generated from a stochastic process according to the following setup [cf. 8]: we fix the number of features $p$, number of datapoints $n$, correlation parameter $\rho=0.3$ and signal to noise parameter $\nu=1$, and generate $\bm{X},\bm{y}$ according to a data generation procedure standard to the literature and stated in Appendix A for brevity.
Following the standard cross-validation paradigm, we then evaluate the cross-validation error for each $\tau$ and $20$ values of $\gamma$ log-uniformly distributed on $[10^{-3},10^{3}]$, using the Generalized Benders Decomposition scheme developed by [7] to solve each MIO to optimality, which is of the form
$\min_{\bm{\beta}\in\mathbb{R}^{p}}\quad$ $\frac{\gamma}{2}\|\bm{\beta}\|_{2}^{2}+\frac{1}{2}\|\bm{X}\bm{\beta}-\bm{y}\|_{2}^{2}\ \text{s.t.}\ \|\bm{\beta}\|_{0}\leq\tau,$ (1)
and selecting the hyperparameter combination with the lowest cross-validation error, for both leave-one-out and five-fold cross-validation.
@@
Figure 1: Leave-one-out (LOOCV, left) and test (right) error for varying $\tau$ and $\gamma$, for an overdetermined setting (top, $n=50,p=10$) and an underdetermined setting (bottom, $n=10,p=50$). In the overdetermined setting, LOOCV is a good estimate of the test error for most values of parameters $(\gamma,\tau)$. In contrast, in the underdetermined setting, LOOCV is a poor approximation of the test error, and estimators that minimize LOOCV ($\gamma\to 0$, $\tau=10$) significantly disappoint out-of-sample. Our conclusions are identical when using five-fold cross-validation (Appendix B).
@@ 1.2 Literature Review
From a statistical learning perspective, there is significant literature on quantifying the out-of-sample performance of models with respect to their training and validation error, originating with the works by [57] on VC-dimension and [11] on algorithmic stability theory. As noted, for instance, by [1], algorithmic stability bounds are generally preferable: they are a posteriori bounds with tight constants that depend on only the problem data. In contrast, VC-dimension bounds are a priori bounds that depend on computationally intractable constants like Rademacher averages. A key conclusion from both streams of work is that simpler and more stable models, as well as models obtained using less computational time to cross-validate, tend to disappoint less out-of-sample. We refer to [26, 52, 3] for more extensive reviews of cross-validation and algorithmic stability.
More recently, the statistical learning theory literature has been connected to the distributionally robust optimization literature by [2, 1, 22, 21, 23] among others. Indeed, [1] proposes solving newsvendor problems by designing decision rules that map features to an order quantity and obtain finite-sample guarantees on out-of-sample costs of newsvendor policies in terms of in-sample cost.
@@
Even closer to our work, [22] proposes correcting solutions to high-dimensional problems by invoking Stein’s lemma to obtain a Stein’s Unbiased Risk Estimator (SURE) approximation of the out-of-sample disappointment and demonstrates that minimizing their bias-corrected training objective generates models that outperform sample-average approximation models out-of-sample. Moreover, they demonstrate that a naive implementation of leave-one-out cross-validation performs poorly in settings with limited data. Building upon this work, [23] proposes debiasing a model’s in-sample performance by incorporating a variance gradient correction term derived via sensitivity analysis. Unfortunately, it is unclear how to extend their approach to the setting considered here, as it applies to problems with linear objectives over subsets of $[0,1]^{n}$.
2 Stability Adjusted Cross-Validation
@@
In this section, we propose techniques for improving the performance of cross-validation. First, we define our notation (Section 2.1) and propose a bound on the test set error of a machine learning model in terms of its $k$-fold error and algorithmic stability, which was originally proven for the special case of leave-one-out cross-validation by [11] (Section 2.2). By leveraging this bound, we propose a technique for improving the performance of cross-validation, namely nested cross-validation with stability regularization (Section 2.3)
@@ 2.1 Setup and Notation
We consider a generic $k$-fold cross-validation setting for supervised learning problems with $n$ datapoints, and our notation is standard to the machine learning literature. Concretely, we have features $(\bm{x}_{1},\ldots,\bm{x}_{n})\in\mathcal{X}^{n}\subseteq\mathbb{R}^{p\times n}$ and response $(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}\subseteq\mathbb{R}^{n}$, and we assume that the data $(\bm{x}_{i},y_{i})_{i\in[n]}\in\mathcal{X}\times\mathcal{Y}$ are drawn i.i.d. from some (unknown) stochastic process. We aim to understand how abstract models $\bm{\beta}:\mathcal{X}\rightarrow\mathcal{Y}$ trained on the full dataset $(\bm{X},\bm{y})$ generalize to other observations from the same stochastic process, by studying $\bm{\beta}$ and related models $\bm{\beta}^{(\mathcal{N}_{j})}:\mathcal{X}\rightarrow\mathcal{Y}$ trained on all data apart from the points $i\in\mathcal{N}_{j}$. We formalize the notion of generalization with a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, e.g., $\ell(\hat{y},y)=(\hat{y}-y)^{2}$. We let $\{\mathcal{N}_{j}\}_{j\in[k]}$ be a partition of the integers $[n]$ into $k$ disjoint subsets, where $k$ is frequently either $5,10$ (five-fold and ten-fold cross-validation) or $n$ (leave-one-out cross-validation), and the cardinality of each fold $\mathcal{N}_{j}$ is typically identical.
To make the dependence of our abstract models $\bm{\beta}$ on hyperparameters explicit, we let $\bm{\theta}\in\Theta$ be the vector of all hyperparameters that are used to select $\bm{\beta}$, and $\Theta$ denote the set of all possible hyperparameter combinations. We define an abstract model trained using hyperparameters $\bm{\theta}$ by $\bm{\beta}(\bm{\theta})$ for concreteness and assume that this choice is unique for simplicity [^1] Relaxing this assumption leads to a notion of Pareto Cross-Validation analogous to Pareto-optimal Benders cuts as described by [37] and Pareto robust optimization as described by [33]. We leave this for future work.. Finally, at the risk of overloading notation, we let, $\bm{\beta}(\bm{\theta},\bm{x}_{i})\in\mathcal{Y}$ denote the output predicted by model $\bm{\beta}(\bm{\theta})$ with data $\bm{x}_{i}$.
Given the above notation, the $k$-fold cross-validation error with hyperparameters $\bm{\theta}$ is given by [^2] We could also consider average over all ($n\choose k$ folds of size $k$ for accuracy. We leave this for future work.:
@@
$h_{j}(\bm{\theta}):=\sum_{i\in\mathcal{N}_{j}}\ell\left(y_{i},\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i})\right).$ (3)
@@
We now define the stability of our models, following [11]. Specifically, let $\mu_{h}$ be the hypothesis stability of our learner analogously to [11, Definition 3] but where $k<n$ folds are possible:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu_{h}:=\max_{j\in[k]}\mathbb{E}_{\bm{x}_{i},y_{i}}\left|\ell\left(y_{i},\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i})\right)-\ell\left(y_{i},\beta(\bm{\theta},\bm{x}_{i})\right)\right|,$ (4)
where the expectation is taken over all $(\bm{x}_{i},y_{i})$ drawn i.i.d. from the underlying stochastic process that generated the training set. This quantity measures the worst-case average absolute change in the loss after omitting a fold of data.
Unfortunately, the hypothesis stability involves computing a possibly high-dimensional expectation and thus is ##\##P-hard to compute in general (even when $\bm{\beta}(\bm{\theta})$ is computable in polynomial time) by reduction from two-stage stochastic programming with random recourse [24, 6]. Moreover, the stochastic process which $(\bm{x}_{i},y_{i})$ is drawn from is often unknown in practice, and thus cannot be computed even with the help of Monte-Carlo simulation or similar techniques to evaluate high-dimensional integrals. Indeed, we have the following result:
Proposition 1.
Determining the quantity $\mu_{h}$ in (4) is ##\##P-hard, i.e., at least as hard as computing the number of optimal solutions to an NP-complete problem, even for the simple case of a binary loss function and i.i.d. discrete random data $(\bm{x}_{i},y_{i})$.
@@
We perform a reduction from Bertsimas and Sturt [6, Corollary 2.1], where the authors report that if $X_{1},\ldots,X_{n}$ are i.i.d. discrete random variables with support containing at least $n$ distinct values, then computing $\mathbb{P}(\sum_{i}X_{i}\leq\alpha)$ is ##\##P-hard. Specifically, let the problem data and lower level training problem be such that
$\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i}))=1\left\{\sum_{i\in[n]\backslash\mathcal{N}_{j}}\sum_{l\in[p]}x_{i,j}\leq\alpha\right\}$
and
$\ell(y_{i},\bm{\beta}(\bm{\theta},\bm{x}_{i}))=1\left\{\sum_{i\in[n]}\sum_{l\in[p]}x_{i,j}\leq\alpha\right\}.$
Thus, if $\mu_{h}$ were not ##\##P-hard then $\mathbb{P}(\sum_{i=1}^{n}X_{i}\leq\alpha)$ would also not be ##\##P-hard, a contradiction. ∎
Thus, to avoid the cost of computing the potentially unknowable and ##\##P-hard quantity $\mu_{h}$, in our computations we approximate (4) via the empirical hypothesis stability analogously to [11, Definition 3] but where $k<n$ folds are possible:
@@
This can be viewed as a sample-average approximation of the hypothesis stability [see 49, for a general theory], and thus it converges almost surely to the true hypothesis stability as $n\rightarrow\infty$ under the usual assumptions of the sample-average-approximation method [49].
We remark that the above empirical hypothesis stability differs from the pointwise hypothesis stability defined by [11, Definition 4] as we average over all data points, rather than only data points omitted when training $\bm{\beta}^{(\mathcal{N}_{j})}$. This is because the pointwise stability only gives generalization bounds on the training error [11], while we are interested in bounds on the kCV error.

### v2

@@ Abstract
We revisit the problem of ensuring strong test set performance via cross-validation, and propose a nested k-fold cross-validation scheme that selects hyperparameters by minimizing a weighted sum of the usual cross-validation metric and an empirical model-stability measure. The weight on the stability term is itself chosen via a nested cross-validation procedure. This reduces the risk of strong validation set performance and poor test set performance due to instability. We benchmark our procedure on a suite of $13$ real-world datasets, and find that, compared to $k$-fold cross-validation over the same hyperparameters, it improves the out-of-sample MSE for sparse ridge regression and CART by $4\%$ and $2\%$ respectively on average, but has no impact on XGBoost. It also reduces the user’s out-of-sample disappointment, sometimes significantly. For instance, for sparse ridge regression, the nested k-fold cross-validation error is on average $0.9\%$ lower than the test set error, while the $k$-fold cross-validation error is $21.8\%$ lower than the test error. Thus, for unstable models such as sparse regression and CART, our approach improves test set performance and reduces out-of-sample disappointment.
Machine Learning, ICML
@@ 1 Introduction
A central problem in machine learning involves constructing models that reliably generalize well to unseen data. One of the most popular approaches is cross-validation as introduced by (Stone, 1974; Geisser, 1975), which selects hyperparameters that perform well on a cross-validation set as a proxy for strong test set performance. Moreover, in the statistical learning literature, there is a broad set of conditions under which the cross-validation loss is a good estimator of the test set error (Plutowski et al., 1993; Arlot & Celisse, 2010; Homrighausen & McDonald, 2013; Hastie et al., 2020; Patil et al., 2021), especially when the sample size is large (Kearns & Ron, 1997; Bousquet & Elisseeff, 2002).
@@
However, both theory (e.g. Shao, 1993; Gupta & Rusmevichientong, 2021) and experiments (e.g. Reunanen, 2003; Rao et al., 2008; Stephenson et al., 2021; Bates et al., 2021) have documented an adaptivity gap between validation and test set performance (although not always observed (Roelofs et al., 2019; Bates et al., 2021)), especially in settings with limited data. Concretely, an adaptivity gap arises when the validation error is systematically lower than the test set error of an ML model.
@@
One explanation for adaptivity gaps is as follows: validation sets yield approximately unbiased estimates of test set performance for a fixed combination of hyperparameters. However, validation scores are random variables and subject to some variance. Therefore, the act of optimizing the validation set error risks selecting hyperparameter combinations that disappoint out-of-sample. In the extreme case where the number of hyperparameters is much larger than the number of samples, optimizing the (cross) validation error can be viewed as training on the (cross) validation set. This phenomenon is well documented in different parts of the statistics and optimization literature, where it is variously called “post-decision surprise”, “out-of-sample disappointment”, “researcher degrees of freedom” or “the optimizer’s curse” (Harrison & March, 1984; Ng, 1997; Smith & Winkler, 2006; Simmons et al., 2011; Van Parys et al., 2021). This raises the question: is it possible to improve the performance of (cross) validation by combating the adaptivity gap, possibly by selecting models with a higher validation error and a lower score according to a second metric.
@@
In this work, we propose a strategy for mitigating out-of-sample disappointment and show that it sometimes improves cross-validation’s performance. Specifically, we propose selecting models according to a weighted sum of their cross-validation error and their empirical hypothesis stability (see Section 2.2). This is motivated by the observation that both the cross-validation error and the hypothesis stability appear in generalization bounds for test set error; thus, minimizing the cross-validation error alone may lead to the selection of high-variance models that perform poorly out of sample.
@@
Our main contributions are threefold. First, we extend a generalization bound on the test set error due to (Bousquet & Elisseeff, 2002) from leave-one-out to $k$-fold cross-validation (kCV). This generalization bound takes the form of the cross-validation error plus a term related to a model’s hypothesis stability. Second, motivated by this (often conservative) bound, we propose regularizing cross-validation by selecting models that minimize a weighted sum of a validation metric and the hypothesis stability, rather than the validation score alone, to mitigate out-of-sample disappointment without being overly conservative. Indeed, models with a low cross-validation error that are stable generalize better than models with a low cross-validation error that are unstable. Moreover, to select the weight in this scheme, we embed the entire scheme within a nested cross-validation procedure. Finally, we empirically evaluate our proposal using sparse ridge regression, CART, and XGBoost, and find that it improves the out-of-sample performance of sparse ridge regression and CART by $4\%$ on average, but has no impact on XGBoost, likely because XGBoost is more stable.
@@ 1.1 Motivating Example: Poor Performance of Cross-Validation for Sparse Linear Regression
We illustrate the pitfalls of cross-validation in a sparse ridge regression setting, as studied by (Bertsimas & Van Parys, 2020; Hazimeh et al., 2022; Liu et al., 2023). Suppose we wish to recover a $\tau_{\text{true}}=5$ sparse regressor which is generated from a stochastic process according to the following setup (cf. Bertsimas et al., 2020): we fix the number of features $p$, number of data points $n$, correlation parameter $\rho=0.3$ and signal to noise parameter $\nu=1$, and generate $\bm{X},\bm{y}$ according to a data generation procedure standard to the literature and stated in Appendix A for brevity.
Following the standard cross-validation paradigm, we then evaluate the cross-validation error for each $\tau$ and $20$ values of $\gamma$ log-uniformly distributed on $10^{-3},10^{3}]$, using the Generalized Benders Decomposition scheme developed by (Bertsimas & Van Parys, 2020) to solve each MIO to optimality, which is of the form
$\min_{\bm{\beta}\in\mathbb{R}^{p}}\quad$ $\frac{1}{2\gamma}\|\bm{\beta}\|_{2}^{2}+\frac{1}{2}\|\bm{X}\bm{\beta}-\bm{y}\|_{2}^{2}\ \text{s.t.}\ \|\bm{\beta}\|_{0}\leq\tau,$ (1)
and selecting the hyperparameter combination with the lowest leave-one-out cross-validation error.
@@
Figure 1: Leave-one-out (LOOCV, left) and test (right) error for varying $\tau$ and $\gamma$, for an overdetermined setting (top, $n=50,p=10$) and an underdetermined setting (bottom, $n=10,p=50$). In the overdetermined setting, LOOCV is a good estimate of the test error for most values of parameters $(\gamma,\tau)$. In contrast, in the underdetermined setting, LOOCV is a poor approximation of the test error, and estimators that minimize LOOCV ($\gamma=10^{3}$, $\tau=10$) significantly disappoint out-of-sample. Our conclusions are identical when using five-fold cross-validation (Appendix B).
@@ 1.2 Literature Review
From a statistical learning perspective, there is significant literature on quantifying the out-of-sample performance of models with respect to their training and validation error, originating with the works by (Vapnik, 1999) on VC-dimension and (Bousquet & Elisseeff, 2002) on algorithmic stability theory. As noted, for instance, by Ban & Rudin (2019), algorithmic stability bounds are generally preferable: they are a posteriori bounds with tight constants that depend on only the problem data. In contrast, VC-dimension bounds are a priori bounds that depend on computationally intractable constants. A key conclusion from both streams of work is that more stable models tend to disappoint less out-of-sample. We refer to Hardt et al. (2016); Stephenson et al. (2021); Bates et al. (2021) for reviews of cross-validation and algorithmic stability.
Relatedly, Kale et al. (2011) introduces a notion of mean-square stability, and shows that, for learning algorithms satisfying this stability property, the variance of the $k$-fold cross-validation error can drop dramatically relative to a single hold-out estimate. Similarly, Lim & Yu (2016) proposes Estimation Stability with Cross-Validation (ESCV) for tuning the Lasso, namely, augmenting cross-validation with an estimation-stability criterion that tends to select smaller and more stable sparse models for the Lasso. Further, Meinshausen & Bühlmann (2010) define stability differently from this paper—in terms of the probability that a variable is selected in a statistical model—and apply their definition to variable selection in regression. All three works are complementary: rather than using stability to justify CV’s accuracy or to stabilize Lasso model selection, we regularize the CV objective with a hypothesis-stability penalty and tune its weight via nested cross-validation.
@@
More recently, the statistical learning theory literature has been connected to distributionally robust optimization by (Ban & Rudin, 2019; Gupta & Rusmevichientong, 2021; Gupta et al., 2024). Indeed, Ban & Rudin (2019) proposes solving newsvendor problems by designing decision rules that map features to an order quantity and obtain finite-sample guarantees on out-of-sample costs of newsvendor policies. Closer to our work, Gupta & Rusmevichientong (2021) proposes correcting solutions to high-dimensional problems by invoking Stein’s lemma to obtain a Stein’s Unbiased Risk Estimator (SURE) approximation of the out-of-sample disappointment. Moreover, they demonstrate that a naive implementation of leave-one-out cross-validation performs poorly with limited data. Building upon this, Gupta et al. (2024) de-bias a model’s in-sample performance by incorporating a variance gradient correction term derived via sensitivity analysis. Unfortunately, it is unclear how to extend their approach to our setting, as it applies to problems with linear objectives over subsets of $0,1]^{n}$.
2 Stability Regularized Cross-Validation
@@
In this section, we propose techniques for improving the performance of cross-validation. First, we define our notation (Section 2.1) and propose a bound on the test set error of a machine learning model in terms of its $k$-fold error and algorithmic stability, which was originally proven for the special case of leave-one-out cross-validation by (Bousquet & Elisseeff, 2002) (Section 2.2). By leveraging this bound, we propose a technique for improving the performance of cross-validation, namely nested cross-validation with stability regularization (Section 2.3)
@@ 2.1 Setup and Notation
We consider a generic $k$-fold cross-validation setting for supervised learning problems with $n$ datapoints, and our notation is standard to the machine learning literature. Concretely, we have features $(\bm{x}_{1},\ldots,\bm{x}_{n})\in\mathcal{X}^{n}\subseteq\mathbb{R}^{n\times p}$ and response $(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}\subseteq\mathbb{R}^{n}$, and we assume that the data $(\bm{x}_{i},y_{i})_{i\in[n]}\in\mathcal{X}\times\mathcal{Y}$ are drawn i.i.d. from some (unknown) stochastic process. We aim to understand how abstract models $\bm{\beta}:\mathcal{X}\rightarrow\mathcal{Y}$ trained on the full dataset $(\bm{X},\bm{y})$ generalize to other observations from the same stochastic process, by studying $\bm{\beta}$ and related models $\bm{\beta}^{(\mathcal{N}_{j})}:\mathcal{X}\rightarrow\mathcal{Y}$ trained on all data apart from the points $i\in\mathcal{N}_{j}$. We formalize the notion of generalization with a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, e.g., $\ell(y,\hat{y})=(\hat{y}-y)^{2}$. We let $\{\mathcal{N}_{j}\}_{j\in[k]}$ be a partition of the integers $[n]$ into $k$ disjoint subsets, where $k$ is frequently either $5,10$ (five-fold and ten-fold cross-validation) or $n$ (leave-one-out cross-validation), and the cardinality of each fold $\mathcal{N}_{j}$ is typically identical.
To make the dependence of our abstract models $\bm{\beta}$ on hyperparameters explicit, we let $\bm{\theta}\in\Theta$ be the vector of all hyperparameters that are used to select $\bm{\beta}$, and $\Theta$ denote the set of all possible hyperparameter combinations. We define an abstract model trained using hyperparameters $\bm{\theta}$ by $\bm{\beta}(\bm{\theta})$ for concreteness and assume that this choice is unique for simplicity. Finally, at the risk of overloading notation, we let $\bm{\beta}(\bm{\theta},\bm{x}_{i})\in\mathcal{Y}$ denote the output predicted by model $\bm{\beta}(\bm{\theta})$ with data $\bm{x}_{i}$.
Given the above notation, the $k$-fold cross-validation error with hyperparameters $\bm{\theta}$ is given by:
@@
$h_{j}(\bm{\theta})=\sum_{i\in\mathcal{N}_{j}}\ell\left(y_{i},\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i})\right).$ (3)
@@
We now define the stability of our models, following (Bousquet & Elisseeff, 2002). Specifically, let $\mu_{h}$ be the hypothesis stability of our learner analogously to (Bousquet & Elisseeff, 2002, Definition 3) but where $k<n$ folds are possible:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu_{h}:=\max_{j\in[k]}\mathbb{E}_{\mathcal{S}\sim\mathcal{D}}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}$ $\bigg|\ell\left(y,\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x})\right)$ (4)
$-\ell\bigg(y,\beta(\bm{\theta},\bm{x})\bigg)\bigg|,$
i.e., the expectation over the training set $\mathcal{S}$ sampled from an underlying stochastic process $\mathcal{D}$, and any relevant fold split randomness, evaluated at a point $(\bm{x},y)\sim\mathcal{D}$ not necessarily contained in the training set. This quantity measures the worst-case average absolute change in the loss after omitting a fold of data.
Unfortunately, the hypothesis stability involves computing a possibly high-dimensional expectation and thus is #\#P-hard to compute in general (even when $\bm{\beta}(\bm{\theta})$ is computable in polynomial time) by reduction from two-stage stochastic programming with random recourse (Hanasusanto et al., 2016; Bertsimas & Sturt, 2020). Moreover, the stochastic process from which $(\bm{x}_{i},y_{i})$ is drawn is often unknown in practice, and thus cannot be computed even with the help of Monte-Carlo simulation or similar techniques to evaluate high-dimensional integrals. We have the following result:
Proposition 2.1.
Determining the quantity $\mu_{h}$ in (4) is #\#P-hard, i.e., at least as hard as computing the number of optimal solutions to an NP-complete problem, even for the simple case of a binary loss function and i.i.d. discrete random data $(\bm{x}_{i},y_{i})$.
@@
We perform a reduction from Bertsimas & Sturt (2020, Corollary 2.1), where the authors report that if $X_{1},\ldots,X_{m}$ are i.i.d. discrete random variables with support containing at least $m$ distinct values, then computing $\mathbb{P}(\sum_{i}X_{i}\leq\alpha)$ is #\#P-hard.
[…]
$(A(T))(x)\;:=\;\begin{cases}\mathbf{1}\{\sum_{z\in T}x(z)\leq\alpha\},&|T|=m,\\ 0,&|T|=2m.\end{cases}$
This algorithm runs in time $O(|T|)$.
Now draw a sample $S=(z_{1},\ldots,z_{2m})\sim\mathcal{D}^{2m}$ and let $S^{(-j)}:=(z_{i})_{i\notin N_{j}}$ denote the sample with fold $j$ removed. By construction, $A(S)=0$, while for each $j\in\{1,2\}$, $A(S^{(-j)})=\mathbf{1}\{\sum_{i\notin N_{j}}x_{i}\leq\alpha\}$, where $x_{i}$ is the feature in $z_{i}$. Since the predictors are constant, the inner expectation over an independent $z\sim\mathcal{D}$ is immaterial and
$\mu_{h}=$ $\max_{j\in\{1,2\}}\mathbb{E}_{S\sim\mathcal{D}^{2m}}\mathbb{E}_{z\sim\mathcal{D}}\Bigl[\bigl|\ell(z,A(S^{(-j)}))-\ell(z,A(S))\bigr|\Bigr]$
$=\mathbb{P}\!\left(\sum_{i=1}^{m}X_{i}\leq\alpha\right)=p.\qed$
Thus, to avoid the cost of computing the potentially unknowable and #\#P-hard quantity $\mu_{h}$, we propose approximating (4) via the empirical hypothesis stability analogously to (Bousquet & Elisseeff, 2002, Definition 3) but where $k<n$ folds are possible:
@@
This can be viewed as a sample-average approximation of the hypothesis stability (see Shapiro et al., 2021, for a general theory), and thus it converges almost surely to the true hypothesis stability as $n\rightarrow\infty$ under the usual assumptions of the sample-average-approximation method (Shapiro et al., 2021).
We remark that the above empirical hypothesis stability differs from the pointwise hypothesis stability defined by (Bousquet & Elisseeff, 2002, Definition 4) as we average over all data points, rather than only data points omitted when training $\bm{\beta}^{(\mathcal{N}_{j})}$. This is because the pointwise stability only gives generalization bounds on the training error (Bousquet & Elisseeff, 2002), while we are interested in bounds on the kCV error.